OpenStudy (anonymous):

Serena saved forty percent of the money she made babysitting last week. If she made sixty-five dollars last week, how much did she save? Show your work.

OpenStudy (anonymous):

@Purplerainbowcherry

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@IMStuck

OpenStudy (kropot72):

One percent of $65 is: \[\large \frac{$65}{100}\] So to find 40% we multiply the value of one percent by 40 giving: \[\large \frac{65}{100} \times\frac{40}{1}=\frac{65\times40}{100}=$\ (you\ can\ calculate)\]

OpenStudy (anonymous):

yeah I can and it's 26?

OpenStudy (kropot72):

You are correct!

OpenStudy (anonymous):

:D THANK YOU!!

OpenStudy (kropot72):

You're welcome :)